40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 23:55:51 2019, gautam, Update, ALS, ALS delay line electronics delayLineDemod.pdf.zip
    Reply  Thu Mar 7 01:06:38 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Tue Mar 12 22:51:25 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Wed Mar 13 01:27:30 2019, gautam, Update, ALS, ALS delay line electronics DFDcal.pdfDFDnoise.pdf
          Reply  Sun Mar 17 00:42:20 2019, gautam, Update, ALS, NF1611 cannot be shot-noise limited? PDnoise.pdf
             Reply  Thu Mar 28 19:40:02 2019, gautam, Update, ALS, BeatMouth with NF1611s assembled IMG_7381.JPGIMG_7382.JPGrelTF_schem.pdfdarkNoise.pdf
                Reply  Fri Mar 29 21:00:06 2019, gautam, Update, ALS, BeatMouth with NF1611s installed IMG_7384.JPGIMG_7385.JPG
                   Reply  Sun Mar 31 15:05:53 2019, gautam, Update, ALS, Fiber beam-splitters not PM IMG_7384.JPG
                      Reply  Wed Apr 3 09:04:01 2019, gautam, Update, ALS, Note about new fiber couplers 
                         Reply  Wed Apr 3 12:32:33 2019, Koji, Update, ALS, Note about new fiber couplers 
                      Reply  Fri Apr 5 00:33:58 2019, gautam, Update, ALS, Promising IR ALS noise IMG_7388.JPGfreeSpace_IR_beat.pdf
                         Reply  Fri Apr 5 11:49:30 2019, gautam, Update, ALS, PSL + X green beat recovery 
                         Reply  Mon Apr 8 00:04:08 2019, gautam, Update, ALS, IR ALS noise budget DFDnoise.pdf
                            Reply  Mon Apr 8 18:28:25 2019, gautam, Update, ALS, EX Green PDH checkout IMG_7393.JPGgreenModeMatch.pdf
                               Reply  Mon Apr 8 23:52:09 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated errMonCalib.pdferrMon.pngOLTF.pdfEX_frequencyNoises.pdf
                                  Reply  Tue Apr 9 00:16:22 2019, rana, Update, ALS, EX Green PDH error monitor calibrated 
                                     Reply  Tue Apr 9 00:18:19 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated 
                                  Reply  Tue Apr 9 18:44:00 2019, gautam, Update, ALS, EX Green PDH discriminant measurement cavityFlashes.pdfmodelPDH.pdf
                            Reply  Tue Apr 9 19:07:12 2019, gautam, Update, ALS, IR ALS noise budget ALS_noiseBudget.pdfALS_noiseBudget.pdf
                               Reply  Thu Apr 11 01:10:05 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X EX_PDH_2kNoise.pdf
                                  Reply  Wed Apr 17 00:50:17 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDHnoise.pdf
                                     Reply  Wed Apr 17 11:01:49 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDH_specGram.pdf
                      Reply  Wed May 29 18:13:25 2019, gautam, Update, ALS, Fiber beam-splitters are now PM 
                         Reply  Fri May 31 15:55:16 2019, gautam, Update, ALS, PSL + X beat restored 
Message ID: 14510     Entry time: Wed Apr 3 09:04:01 2019     In reply to: 14503     Reply to this: 14513
Author: gautam 
Type: Update 
Category: ALS 
Subject: Note about new fiber couplers 

The new fiber beam splitters we are ordering, PFC-64-2-50-L-P-7-2-FB-0.3W, have the slow axis working and fast axis blocked. The way the light is coupled into the fibers right now is done to maximize the amount of light into the fast axis. So we will have to do a 90deg rotation if we use that part. Probably the easiest thing to do is to put a HWP immediately before the free-space-to-fiber collimator.

Update 6pm: They have an "SB" version of the part with the slow axis blocked and fast axis enabled, same price, so I'll ask Chub to get it.

ELOG V3.1.3-